### v1

@@ 4.3 Latent Space Analysis
For the registration approach, we analyzed the latent priors and their relation to the deformation. We again performed a PCA on the latent priors and relate the first two principal components (explaining around 13 % of the variance) with the mean Jacobian determinant that is computed using the retina as a mask. Fig. 11 shows the results, indicating a significant correlation between the mean Jacobian determinant – which quantifies the volume changes caused by the deformation – and the second principal component of the latent space.
@@ 5 Discussion
Our proposed interpolation method achieves better performance than the considered competitive approaches. For healthy cases, it can interpolate the retinal layer shape most accurately while using even larger slice distances than clinical OCT volumes. On pathological data, we were able to repeat good OCT B-scan interpolation performance while retaining pathologies within these interpolated B-scans. Our approach is mainly limited by the low-frequency reconstruction by the generalizable INR, leaving a disparity in intensity reconstruction. Despite that, the generalizable INR enables the incorporation of en-face modality information and segmentation prediction of unseen OCT volumes. This in turn results in better reconstruction of retina shape, vessel structures, and fluid positions in interpolated slices. Further improvement might be gained by incorporating sub-pixel-level layer boundaries instead of the voxel-wise segmentation used in this work. Shape-based interpolation methods similarly make use of segmentations and boundaries and are well-suited for interpolating label maps, but do not address the B-scan interpolation task because intensity information is not modeled. In contrast, our approach generates supersampled OCT data with corresponding labels, which can benefit various downstream applications, such as volumetric segmentation, 3D reconstruction, longitudinal analyses or image registration as in Sec. 4.2. Demonstrating direct clinical benefit remains an important direction for future work and will be addressed in subsequent studies.
@@
To provide further analysis and understanding of the population, we built upon the previous findings and proposed a similar framework for registration and atlas generation. We combine methods from grossbrohmer2024sina and wolterink2022implicit with a generalization approach to obtain a pipeline that enables the joint prediction of a resolution-agnostic atlas and corresponding deformation fields. grossbrohmer2024sina proposed an INR as an atlas representation but used a simple registration based on B-spline control points. This instance-based B-spline registration showed good performance on other datasets but – as shown in this study – scales poorly to the challenges associated with OCT. Implicit deformable image registration (wolterink2022implicit) on the other hand, exhibited good alignment performance with close to no foldings on our OCT data. To resolve the restriction of the SingleINR to prior knowledge, we proposed a generalizable INR registration framework to include population-based knowledge and auxiliary information (i.e., retinal layer segmentation) in a joint atlas and deformation optimization process. Thus, we enable the creation of a sharp and accurate atlas representation as well as a continuous representation of the deformation. The resulting method is, by construction, resolution-agnostic and can be used on volumes containing varying numbers of B-scans. Our combined framework improved the alignment and representation performance, reduced the execution time during inference, and reduced the deformation field representation to a single lower-dimensional vector. In line with the findings of tian2024nephi, we show that generalizable INRs present an efficient approach to model the registration process. Since smoothness of the deformation field is intended, the drawback of low-frequency reconstruction observed for interpolation does not transfer to registration.
@@
Altogether, INRs simplify the handling and processing of such high-dimensional image data. Especially for image registration, INRs seem to cause a trend for new approaches (wolterink2022implicit; sideri2024sinr; van2023robust; tian2024nephi; grossbrohmer2024sina). When the image is represented by an INR, it eliminates the need for linear interpolation in the registration process while maintaining differentiability. Furthermore, implicitly representing data as functions reduces memory and runtime requirements when compared to other iterative approaches. This is particularly the case once the INR is combined with a generalization approach and the representation is compressed into the latent priors. With this, new potentials and analysis options arise, as we demonstrated for our proposed methods.
@@ 6 Conclusion
The data is currently not publicly available. The CSCR dataset is acquired in clinical routine and is subject to data protection. The physiological in-house dataset is in preparation to be published and will be referenced in our GitHub repository.

### v3

@@ 4.3 Latent Space Analysis
For the registration approach, we analyzed the latent priors and their relation to the deformation. We again performed a PCA on the latent priors and related the first two principal components (explaining around 13 % of the variance) with the mean Jacobian determinant that is computed using the retina as a mask. Fig. 11 shows the results, indicating a significant correlation between the mean Jacobian determinant – which quantifies the volume changes caused by the deformation – and the second principal component of the latent space.
@@ 5 Discussion
Our proposed interpolation method achieves better performance than the considered competitive approaches. For healthy cases, it can interpolate the retinal layer shape most accurately while using even larger slice distances than clinical OCT volumes. On pathological data, we were able to repeat good OCT B-scan interpolation performance while retaining pathologies within these interpolated B-scans. Our approach is mainly limited by the low-frequency reconstruction by the generalizable INR, leaving a disparity in intensity reconstruction. Despite that, the generalizable INR enables the incorporation of en-face modality information and segmentation prediction of unseen OCT volumes. This in turn results in better reconstruction of retina shape, vessel structures, and fluid positions in interpolated slices. Further improvement might be gained by incorporating sub-pixel-level layer boundaries instead of the voxel-wise segmentation used in this work. Shape-based interpolation methods similarly make use of segmentations and boundaries and are well-suited for interpolating label maps, but do not address the B-scan interpolation task because intensity information is not modeled. In contrast, our approach generates supersampled OCT data with corresponding labels, which can benefit various downstream applications, such as volumetric segmentation, 3D reconstruction, longitudinal analyses or image registration, as in Sec. 4.2. Demonstrating direct clinical benefit remains an important direction for future work and will be addressed in subsequent studies.
@@
To provide further analysis and understanding of the population, we built upon the previous findings and proposed a similar framework for registration and atlas generation. We combine methods from Großbröhmer et al. (2024) and Wolterink et al. (2022) with a generalization approach to obtain a pipeline that enables the joint prediction of a resolution-agnostic atlas and corresponding deformation fields. Großbröhmer et al. (2024) proposed an INR as an atlas representation but used a simple registration based on B-spline control points. This instance-based B-spline registration showed good performance on other datasets but – as shown in this study – scales poorly to the challenges associated with OCT. Implicit deformable image registration (Wolterink et al., 2022) on the other hand, exhibited good alignment performance with close to no foldings on our OCT data. To resolve the restriction of the SingleINR to prior knowledge, we proposed a generalizable INR registration framework to include population-based knowledge and auxiliary information (i.e., retinal layer segmentation) in a joint atlas and deformation optimization process. Thus, we enable the creation of a sharp and accurate atlas representation as well as a continuous representation of the deformation. The resulting method is, by construction, resolution-agnostic and can be used on volumes containing varying numbers of B-scans. Our combined framework improved the alignment and representation performance, reduced the execution time during inference, and reduced the deformation field representation to a single lower-dimensional vector. In line with the findings of Tian et al. (2024), we show that generalizable INRs present an efficient approach to model the registration process. Since smoothness of the deformation field is intended, the drawback of low-frequency reconstruction observed for interpolation does not transfer to registration.
@@
Altogether, INRs simplify the handling and processing of high-dimensional image data. Especially for image registration, INRs seem to cause a trend for new approaches (Wolterink et al., 2022; Sideri-Lampretsa et al., 2024; Van Harten et al., 2023; Tian et al., 2024; Großbröhmer et al., 2024). When the image is represented by an INR, it eliminates the need for linear interpolation in the registration process while maintaining differentiability. Furthermore, implicitly representing data as functions reduces memory and runtime requirements when compared to other iterative approaches. This is particularly the case once the INR is combined with a generalization approach and the representation is compressed into the latent priors. With this, new potentials and analysis options arise, as we demonstrated for our proposed methods.
@@ 6 Conclusion
The CSCR dataset was acquired during clinical routine and is not publicly available due to data protection regulations. The in-house dataset of healthy volunteers is currently being prepared for public release. Further details will be provided in the GitHub repository after release.